OpenStudy (anonymous):

This question has me stumped, please help? Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now? -3 -7 -17 -or 21

OpenStudy (rsadhvika):

Mary is 14 years younger than her sister Jane. => m = j-14 -----------------(1) In four years, Jane will be 3 times older than Mary => j+3 = 3(m+3) -----------------(2)

OpenStudy (anonymous):

Oh...I had the equations set up wrong

OpenStudy (rsadhvika):

hmm happens all the time with me too... lol you can solve the two equations for j and m ?

OpenStudy (anonymous):

yep...Thanks :)

OpenStudy (rsadhvika):

yw :)